OpenStudy (anonymous):

The Us Centers for Disease Control recommends that for a typical 2000- calorie daily diet, no more than 65g of fat can be consumed. In the first 3 days of a 4 day vacation, Teresa consumed 55g, 80g, and 70g of fat. Determine how many grams of fat Teresa can consume on the 4th day if she is to average no more than 65g of fat per day.

Parth (parthkohli):

Average of n number of numbers is: \(\Large \color{MidnightBlue}{\Rightarrow {S_{n} \over n} }\) This means sum of numbers over n.

Parth (parthkohli):

Let's put an equation here, where x = the number of grams the person will eat

OpenStudy (anonymous):

ok

Parth (parthkohli):

\(\Large \color{MidnightBlue}{\Rightarrow {55 + 80 + 70 + x \over 4} = 65 }\)

Parth (parthkohli):

To solve this equation, we must isolate x. This is done my. 1) Multiplying both sides by 4. \(\Large \color{MidnightBlue}{\Rightarrow 55 + 80 + 70 + x = 260 }\) 2) Operating on like terms. \(\Large \color{MidnightBlue}{\Rightarrow 205 + x = 260 }\) 3) Subtracting 205 from both sides. \(\Large \color{MidnightBlue}{\Rightarrow \cancel {205}+x = 260 - 205 = 55 }\)